# Minimal Degrees, Volume Growth, and Curvature Decay on Complete Kähler Manifolds

### v1

@@ 2.1. Structure of Kähler manifolds with maximal volume growth and the tangent cones
Claim: Let $\pi:\widetilde{M}\rightarrow M$ be the covering map, $f\in\mathcal{O}_{P}(M)$, then $\operatorname{deg}(f)=\operatorname{deg}(f\circ\pi)$.
It’s obvious since $\pi(B(\tilde{p},r))=B(p,r)$ for any $\pi(\tilde{p})=p$ and any $r>0$.
@@
Remark 2.8.
Similar argument can also prove the splitting theorem in Section 3 in [7], by using Theorem 4.1 in [23].
2.2. Gradient expanding Kähler-Ricci solitons
Definition 2.9.
@@ 2.2. Gradient expanding Kähler-Ricci solitons
They we list some important properties of gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature which can be found in [2, 6, 30, 4, 28].
Proposition 2.10.
@@
$\frac{\partial}{\partial t}\varphi(t-1)(x)=-\frac{1}{t}\nabla^{g}f(\varphi(t-1)(x)),\varphi(0)=\text{id}.$ (2.6)
@@
$g(t)=t\varphi(t-1)^{*}g$ (2.7)
solves the Kähler-Ricci flow equation on $[0,+\infty)$ with $g(1)=g$.
(5) ((13) in [28]) The potential function $f$ satisfies the estimate
$f(\varphi(O,t))\leq f(O)e^{2t}.$ (2.8)
@@ 3. Proof of Theorem B
The key observation is the following which may be well-known for experts in the fields. However, for reader’s convenience, we give a proof here.
Lemma 3.1.
@@
be the unique Kähler-Ricci flow solution on $[0,+\infty)$ such that $h(1)=h$. Where
@@
Replace $(M,g)$ in Theorem 1.6 by $(N,h(1))$, we conclude that there exists a metric Kähler cone $N_{\infty}$ such that $(N,h(t),O)\rightarrow(N_{\infty},h_{\infty},O_{\infty})$ as $t\rightarrow 0$ in the Gromov-Hausdorff sense. So it remains to prove this cone $N_{\infty}$ is the unique tangent cone of $(N,h(1))$. In fact we have
$(N_{\infty},h_{\infty},O_{\infty})$ $=\lim\limits_{t\to 0}(N,h(t),O)$
$=\lim\limits_{t\to 0}(N,t\varphi(t-1)^{\star}h(1),O)$
$=\lim\limits_{t\to 0}(N,th(1),\varphi(t-1)(O))$
$=\lim\limits_{t\to 0}(N,th(1),(O)).$
All the limits above are in the sense of Gromov-Hausdorff convergence. The second-to-last equality holds by the following claim:
@@
Thus, we now have a clear geometric framework characterizing the long-time asymptotics of the Kähler-Ricci flow on a complete noncompact Kähler manifold with bounded nonnegative bisectional curvature and maximal volume growth, along with its relationship to its tangent cones.
@@
Proof of Theorem B:
Following the notation of Theorem 1.6, take the Poincaré coordinate $(w_{1},\cdots,w_{n})$ as in page 2622 in [28] on $(N,h(1),O)$ and normalize them such that
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B_{h(1)}(O,1)}{w_{i}}\overline{w_{j}}=\delta_{ij},\quad w_{i}(O)=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$
According to the proof of Theorem 1.4 in [28], the aforementioned Poincaré coordinate $(w_{1},\cdots,w_{n})$ is a canonical coordinate on $(N,h(1),O)$. Following the procedure of (5) in Theorem 2.4 on $(N,h(1),O)$ and $(M,g,p)$, we get two canonical coordinates $(z_{1\infty},\cdots,z_{n\infty})$ and $(w_{1\infty},\cdots,w_{n\infty})$ on $N_{\infty}=M_{\infty}$, with degree $\overrightarrow{d_{\operatorname{min}}}(M)$ and $\overrightarrow{d_{\operatorname{min}}}(N)$. Thus from (4) in Theorem 2.4,
@@
Therefore, (1),(2) has been proved. For (3), note that in [18], Liu has proved that $\operatorname{ASCD}(M,g(t))=4n(\sum\limits_{i=1}^{n}d^{(i)}_{\operatorname{min}}-n)=4n(\sum\limits_{i=1}^{n}\mu_{i})$ is invariant along $t$. The proof is complete.
As an application, we can use it to prove Corollary 1.11:
Proof of Corollary 1.11:
Firstly, we assume $(M,g)$ has bounded curvature. Take a canonical coordinate $(f_{1},\cdots,f_{n})$ on $(M,g)$, with degree $\overrightarrow{d_{\operatorname{min}}}$. By comparing the distance function $d_{g(t)}(x,p)$ along the flow using the shrinking balls lemma by Simon and Topping (Corollary 3.3 in [26]), we may conclude that $(f_{1},\cdots,f_{n})$ is still a global coordinate on $(M,g(t))$, with the same degree $\overrightarrow{d_{\operatorname{min}}}$. Let $g_{k}(t)=\frac{1}{t_{i}}g(t_{i}t)$, we normalized them such that
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B_{g_{k}(1)}(O,1)}{f_{ik}}\overline{f_{jk}}=\delta_{ij},\quad f_{ik}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,n\},k\in\mathbb{N}.$
@@
Following the same proof as in (5) in Theorem 2.4, these $(f_{1k},\cdots,f_{nk})$ converges smoothly to $(f_{1\infty},\cdots,f_{n\infty})$ on $(N,h(1))$ and serving as a canonical coordinate on $N$. Due to (3) in Theorem 2.1, WLOG we can assume $(f_{1\infty},\cdots,f_{n\infty})$ is just a Poincaré coordinate on $(N,h(1))$. Therefore
@@
Finally, by the smooth convergence of $g_{i}(1)$ and the proof of Theorem 1.2 in [3], the Corollary is proved. In fact, for any $i$ and any nonzero $v\in V_{1}\oplus\cdots\oplus V_{i}$ but $v\notin V_{1}\oplus\cdots\oplus V_{i-1}$,
@@
always holds.
If $(M,g)$ has unbounded curvature, thanks to Theorem 1.9, above argument can still be applied.
@@ 4. Proof of Theorem A
Firstly we assume $M$ is of nonmaximal volume growth, assume there exists a nonzero polynomial holomorphic $n$-form $s\in H^{0}_{P}(M,\mathcal{K}_{M})$, the Poincaré-Lelong equation states that
@@
where $(s)$ is the zero divisor of $s$. In other words, $h=\log(\|s\|^{2}+1)$ is a smooth plurisubharmonic function of logarithmic growth. Since the universal cover of $M$ does not split, the heat-flow method of Ni-Tam [23] can be applied to perturb $h$ to $h^{\prime}$ which is a smooth plurisubharmonic function of still logarithmic growth (for details, see Theorem 0.1 (iii) in [23]). Then we solve the $\bar{\partial}$-equation by applying the Hörmander $L^{2}$-estimate similarly as the proof of Lemma 3 in [12], there exists a nonconstant polynomial growth holomorphic function on $M$, contradicting with Theorem 2 in [12]. As a result, $D_{\operatorname{min}}=+\infty=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n$.
Now we assume $M$ is of maximal volume growth. Let $(z_{1},\cdots,z_{n})$ be a canonical coordinate of $M$ with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$ as stated in Remark 2.3. By gradient estimate [5],
@@
The proof of the opposite direction is similar to the proof of Corollary 1.5 in [28], but a little more complicated. Let $s\in H^{0}_{D_{\operatorname{min}}}(M,\mathcal{K}_{M})$, then note that Corollary 3 in [13] also holds for holomorphic sections of holomorphic line bundle with nonpoaitive curvature, there exists a constant $C>0$ such that
$M_{s}(r,0)\leq Cr^{D_{\operatorname{min}}}\quad\text{ for any }r>0.$ (4.1)
Fix $t_{k}\rightarrow+\infty$, there exists a long-time Kähler-Ricci flow solution $g(t)$ which satisfies properties in Theorem 1.9. And set $g_{k}(t)=\frac{1}{t_{i}}g(t_{i}t)$, assume $t_{k}\geq 1$.
@@
Proof.
By the shrinking balls lemma by Simon and Topping (Corollary 3.3 in [26]), for any $t_{k}>0$, we have
@@
Also, thank to Corollary 1.10, (1.2) also holds for $M$ with unbounded curvature, which means that
@@
for some nonnegative constant $C$ which relies only on $M$ and is independent of $x$. Also by the Ricci flow equation,
@@
[…]
Thanks to Theorem 1.9, the boundedness condition of curvature in Theorem 1.6 can be removed. Using the notations in Theorem 1.6, recall the sequence $(U_{k},F_{k}^{*}J.F_{k}^{*}(g_{k}(1)),O)$ converges smoothly to $(N,J_{\infty},h(1),O)$ uniformly on compact subsets. Define
$s_{k}=\frac{s\mid_{F_{k}(U_{k})}}{\sup\limits_{B_{g_{k}(1)}(p,1)}\|s\|},\quad\tilde{s}_{k}=F_{k}^{*}s_{k}.$ (4.2)
Since $\mathcal{K}_{M}(g_{k}(1))$ has nonpositive curvature, three-circle theorem in [13] can be applied to get for any fixed $r>0$, for large $k$,
@@
Therefore there exists a constant $C=C(K)$ such that $\|\Omega_{k}\|_{F_{k}^{*}(g_{k}(1))}\geq C(K)$ uniformly on $K\subset N$ compact. Then we can deduces
@@
A much simpler way is consider directly a tangent cone of $M$. Assume $(M,g_{i}=r_{i}^{-2}g)\rightarrow M_{\infty}$ in the Gromov-Hausdorff sense, where $r_{i}\rightarrow\infty$. Take a canonical coordinates $(z_{1},\cdots,z_{n})$ on $(M,g)$, we know that
@@
And from (4) we have known that
@@
Following the notations in Theorem 2.3 (5), and define
@@
As in [18], since $\varphi_{i}$ are psh on $M_{i}$, by passing to a subsequence, we may assume $\varphi_{i}$ converges to a psh function on $M_{\infty}$ in $L_{\operatorname{loc}}^{1}$, denoted by $\varphi=\log|dz|^{2}$. Therefore $e^{\varphi_{i}}\rightarrow e^{\varphi}$ almost everywhere.
By the three circle theorem as in the proof of Theorem 2.3 (5), there exists a $C$ such that

### v2

@@ 2.1. Structure of Kähler manifolds with maximal volume growth and tangent cones
Claim 1.
Let $\pi:\widetilde{M}\rightarrow M$ be the covering map, $f\in\mathcal{O}_{P}(M)$, then $\operatorname{deg}(f)=\operatorname{deg}(f\circ\pi)$.
Proof of Claim 1. It’s obvious since $\pi(B(\tilde{p},r))=B(p,r)$ for any $\pi(\tilde{p})=p$ and any $r>0$. ∎
@@
Remark.
Similar argument can also prove the splitting theorem in Section 3 in [CH25], by using Theorem 4.1 in [NT03].
2.2. Gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature
Definition 3.
@@ 2.2. Gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature
They we list some important properties of gradient expanding Kähler-Ricci solitons with nonnegative Ricci curvature which can be found in [CT05, Bry08, CCGGIIKLLN, Zhang09, CT11, Yang22].
Proposition 1.
@@
(2.11) $\frac{\partial}{\partial t}\varphi(t-1)(x)=-\frac{1}{t}\nabla^{g}f(\varphi(t-1)(x)),\varphi(0)=\text{id}.$
@@
(2.12) $g(t)=t\varphi(t-1)^{*}g$
solves the Kähler-Ricci flow equation on $0,+\infty)$ with $g(1)=g$.
(5) ((13) in [Yang22]) The potential function $f$ satisfies the estimate
(2.13) $f(\varphi(O,t))\leq f(O)e^{2t}.$
(6) There exists a global coordinate called Poincaré coordinates on $N$ which is constructed as following: there exist local coordinates $\left(U_{O},z_{i}\right)$ at $O$ such that $\left\{\frac{\partial}{\partial z_{i}}\right\}$ is unitary at $O$ and the complexified vector field $\nabla f-\sqrt{-1}J\nabla f=\sum\limits_{i=1}^{n}\left(1+\mu_{i}\right)z_{i}\frac{\partial}{\partial z_{i}}$. Then one can extend $\left(U_{O},z_{i}\right)$ to the whole $(N,g)$ by defining
$z_{i}(\psi(p,t))=e^{\left(1+\mu_{i}\right)t}z_{i}(p)$
for any $p\in U_{O}$ where $\psi(p,t)$ is the holomorphic flow generated by $\nabla f$ starting from $p$.
@@ 3. Proof of Theorem B
The key observation is the following which may be well-known for experts in the field. However, for reader’s convenience, we give a proof here.
Lemma 1.
@@
be the unique Kähler-Ricci flow solution on $0,+\infty)$ such that $h(1)=h$. Where
@@
Replace $(M,g)$ in Theorem 2 by $(N,h(1))$, we conclude that there exists a metric Kähler cone $N_{\infty}$ such that $(N,h(t),O)\rightarrow(N_{\infty},h_{\infty},O_{\infty})$ as $t\rightarrow 0$ in the Gromov-Hausdorff sense. So it remains to prove this cone $N_{\infty}$ is the unique tangent cone of $(N,h(1))$.
@@
In fact we have
$(N_{\infty},h_{\infty},O_{\infty})$ $=\lim\limits_{t\to 0}(N,h(t),O)$
$=\lim\limits_{t\to 0}(N,t\varphi(t-1)^{*}h(1),O)$
$=\lim\limits_{t\to 0}(N,th(1),\varphi(t-1)(O))$
$=\lim\limits_{t\to 0}(N,th(1),(O)).$
All the limits above are in the sense of Gromov-Hausdorff convergence. The second-to-last equality holds by Claim 2: ∎
Hence we now have a clear geometric framework characterizing the long-time asymptotics of the Kähler-Ricci flow on a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth, along with its relationship to its tangent cones.
@@
Proof of Theorem B.
Following the notation of Theorem 2, take the Poincaré coordinate $(w_{1},\cdots,w_{n})$ from Proposition 1 on $(N,h(1),O)$ and normalize them such that
$\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B_{h(1)}(O,1)}{w_{i}}\overline{w_{j}}=\delta_{ij},\quad w_{i}(O)=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$
According to the proof of Theorem 1.4 in [Yang22], the aforementioned Poincaré coordinate $(w_{1},\cdots,w_{n})$ is a canonical coordinate on $(N,h(1),O)$. Following the procedure of (5) in Theorem 5 on $(N,h(1),O)$ and $(M,g,p)$, we get two canonical coordinates $(z_{1\infty},\cdots,z_{n\infty})$ and $(w_{1\infty},\cdots,w_{n\infty})$ on $N_{\infty}=M_{\infty}$, with degree $\overrightarrow{d_{\operatorname{min}}}(M)$ and $\overrightarrow{d_{\operatorname{min}}}(N)$. Thus from (4) in Theorem 5,
@@
Therefore, (1),(2) has been proved. For (3), note that in [Liu24], Liu has proved that $\operatorname{ASCD}(M,g(t))=4n(\sum\limits_{i=1}^{n}d^{(i)}_{\operatorname{min}}-n)=4n(\sum\limits_{i=1}^{n}\mu_{i})$ is invariant along $t$. The proof is complete. ∎
As an application, we can use it to prove Corollary 1:
Proof of Corollary 1:
Take a canonical coordinate $(f_{1},\cdots,f_{n})$ on $(M,g)$, with degree $\overrightarrow{d_{\operatorname{min}}}$. By comparing the distance function $d_{g(t)}(x,p)$ along the flow using the shrinking balls lemma by Simon and Topping (Corollary 3.3 in [ST22]), we may conclude that $(f_{1},\cdots,f_{n})$ is still a global coordinate on $(M,g(t))$, with the same degree $\overrightarrow{d_{\operatorname{min}}}$. Let $g_{k}(t)=\frac{1}{t_{i}}g(t_{i}t)$, we normalize them such that
$\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B_{g_{k}(1)}(O,1)}{f_{ik}}\overline{f_{jk}}=\delta_{ij},\quad f_{ik}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,n\},k\in\mathbb{N}.$
@@
Following the similar proof as in (5) in Theorem 5, these $(f_{1k},\cdots,f_{nk})$ converges smoothly to $(f_{1\infty},\cdots,f_{n\infty})$ on $(N,h(1))$ and serves as a canonical coordinate on $N$. Due to (3) in Theorem 4, we can assume $(f_{1\infty},\cdots,f_{n\infty})$ is just a Poincaré coordinate on $(N,h(1))$. Therefore
@@
Finally, by the smooth convergence of $g_{i}(1)$ and the proof of Theorem 1.2 in [CT06], the Corollary is proved. In fact, for any $i$ and any nonzero $v\in V_{1}\oplus\cdots\oplus V_{i}$ but $v\notin V_{1}\oplus\cdots\oplus V_{i-1}$,
@@
always holds. ∎
@@ 4. Proof of Theorem A
Firstly we assume $M$ is of nonmaximal volume growth, assume there exists a nonzero polynomial growth holomorphic $n$-form $s\in H^{0}_{P}(M,\mathcal{K}_{M})$, the Poincaré-Lelong equation states that
@@
where $(s)$ is the zero divisor of $s$. In other words, $h:=\log(\|s\|^{2}+1)$ is a smooth plurisubharmonic function of logarithmic growth. Since the universal cover of $M$ does not split, the heat-flow method of Ni-Tam [NT03] can be applied to perturb $h$ to $h^{\prime}$ which is a smooth plurisubharmonic function of still logarithmic growth. Then we solve the $\bar{\partial}$-equation by applying the Hörmander $L^{2}$-estimate to obtain a nonconstant polynomial growth holomorphic function on $M$, contradicting with Theorem 1. As a result, $D_{\operatorname{min}}=+\infty=\sum\limits_{i=1}^{n}d_{\operatorname{min}}^{(i)}-n$.
Now we assume $M$ is of maximal volume growth. Let $(z_{1},\cdots,z_{n})$ be a canonical coordinate of $M$ with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$. By gradient estimate [CY75],
@@
We will give two distinct proofs of the opposite direction:
Proof I.
The first proof of the opposite direction is similar to the proof of Corollary 1.5 in [Yang22], but a little more complicated. Let $s\in P(M,\mathcal{K}_{M})$ with degree $D_{\operatorname{min}}$, then note that Corollary 3 in [Liu16c] can also apply on holomorphic sections of holomorphic line bundle with nonpositive curvature, there exists a constant $C>0$ such that
(4.14) $M_{s}(r,0)\leq Cr^{D_{\operatorname{min}}}\quad\text{ for any }r>0.$
Fix $t_{k}\rightarrow+\infty$, there exists a long-time Kähler-Ricci flow solution $g(t)$ which satisfies properties in Theorem 2. And set $g_{k}(t)=\frac{1}{t_{i}}g(t_{i}t)$, assume $t_{k}\geq 1$.
@@
Proof of Claim 3.
By the shrinking balls lemma by Simon and Topping (Corollary 3.3 in [ST22]), for any $t_{k}>0$, we have
@@
Then by (1.2) we obtain that
@@
for some nonnegative constant $C$ which relies only on $M$ and is independent of $x$. From Ricci flow equation,
@@
Using the notations in Theorem 2, recall the sequence $(U_{k},F_{k}^{*}J.F_{k}^{*}(g_{k}(1)),O)$ converges smoothly to $(N,J_{\infty},h(1),O)$ uniformly on compact subsets. Define
(4.15) $s_{k}=\frac{s\mid_{F_{k}(U_{k})}}{\sup\limits_{B_{g_{k}(1)}(p,1)}\|s\|},\quad\tilde{s}_{k}=F_{k}^{*}s_{k}.$
Since $\mathcal{K}_{M}(g_{k}(1))$ has nonpositive curvature, three-circle theorem can be applied to get for any fixed $r>0$, for large $k$,
@@
Therefore there exists a constant $C=C(K)$ such that $\|\Omega_{k}\|_{F_{k}^{*}(g_{k}(1))}\geq C(K)$ uniformly on $K\subset N$ compact. Then we can deduce
@@
Proof II.
A much simpler way is considering directly a tangent cone of $M$. Assume $(M,g_{i}=r_{i}^{-2}g)\rightarrow M_{\infty}$ in the Gromov-Hausdorff sense, where $r_{i}\rightarrow\infty$. Take a canonical coordinates $(z_{1},\cdots,z_{n})$ on $(M,g)$, we know that
@@
We have known that
@@
Following the notations in Theorem 5 (5), and define
@@
As in [Liu24], since $\varphi_{i}$ is plurisubharmonic on $M_{i}$, by passing to a subsequence, we may assume $\varphi_{i}$ converges to a plurisubharmonic function on $M_{\infty}$ in $L_{\operatorname{loc}}^{1}$, denoted by $\varphi=\log|dz|^{2}$. Therefore $e^{\varphi_{i}}\rightarrow e^{\varphi}$ almost everywhere.
By the three circle theorem as in the proof of Theorem 5 (5), there exists a $C$ such that